OpenStudy (anonymous):

What is the sum of the first 14 terms in the Fibonacci Sequence? Also what is the 14th term

OpenStudy (anonymous):

anybody?

OpenStudy (baldymcgee6):

F0 F1 F2 F3 F4 F5 F6 F7 F8 F9 F10 F11 F12 F13 F14 0 1 1 2 3 5 8 13 21 34 55 89 144 233 377

OpenStudy (anonymous):

whats the sum?

OpenStudy (anonymous):

0+1+1+2+3+5+8+13+21+34+55+89+144+233+377=986

OpenStudy (anonymous):

Must be nice to have people do all your work for you.

OpenStudy (calculusfunctions):

A Fibonacci sequence is defined recursively as\[F _{n +2}=F _{n}+F _{n +1}\]with\[F _{1}=F _{2}=1\]and n is any natural number greater than or equal to 1.

OpenStudy (anonymous):

thanks you all! I get it now